OpenStudy (anonymous):

A bank recommends that the maximum monthly payment for a home be 1/3 of your total monthly income. If your monthly income is $3780. What would the bank recommend as your maximum monthly house payment? $ ? The bank estimates that 1/5 of your monthly income remaining AFTER you have paid your mortgage should be budgeted for housing-related costs (e.g. property tax, heat, hydro, water, etc.). How much should you budget each month for housing-related costs? $ ?

OpenStudy (texaschic101):

1/3 of 3780 = 1/3 * 3780 = 3780/3 = $ 1260 (maximum monthly house payment) 3780 - 1260 = 2520 (money left after paying mortgage) 1/5 of 2520 = 1/5 * 2520 = 2520/5 = $ 504 (budgeted each month for housing related costs)

OpenStudy (anonymous):

thank you again :)

OpenStudy (texaschic101):

anytime :)